Determine whether this function f(n) = n2 − 1 is one-to-one, onto, or both. Explain your answers. The domain of each function is the set of all integers. The codomain of each function is also the set of all integers


1
Expert's answer
2021-10-26T03:17:43-0400

Let us determine whether the function "f:\Z\to\Z,\ f(n) = n^2 − 1," is one-to-one, onto, or both.

Since "f(-1)=(-1)^2-1=0=1^2-1=f(1)," we conclude that this function is not one-to-one. Taking into account that for "y=-2" the equation "f(n)=y," that is "n^2-1=-2" or "n^2=-1," has no integer solutions, we conclude that "f^{-1}(-2)=\emptyset," and hence the function "f" is not onto. Therefore, "f" is not a bijection.